a loop , A straight wire carries a 6.0A current that is directed in a positive x direction. When a uniform magnetic fileld is applied perpendicular to a 3.0 m segment of a site. the magnetic force on the segment is 0.36 N, directed in the negative y direction . what is the magnitude and direction of the magnetic field?
1
Expert's answer
2020-10-12T07:47:15-0400

Since the ampere force (magnetic force) is equal to "F_A=B\times I\times l\times\sin\alpha" , where "B" is magnetic induction (magnitude characterizing magnetic field), "I" is current strength, "l" is the length of the conductor, "\alpha" is the angle between vectors "\vec{I}" and "\vec B" , therefore

"B =\frac {F_A}{I\times l\times\sin\alpha}" "=\frac {0.36}{6\times 3 \times\sin (90\degree)}=0.02" T.

According to the law of the left hand, the direction of magnetic induction will be directed in the negative direction of the "Z" axis.